AFWEX Status, 03 December 2000

Weather was iffy again today, and we scheduled for a 1745 CST DC-8 take-off time with up to a 4 hour delay. After dawn, there appeared to be an opportunity for clear skies at the CART site. The DC-8 then proceeded to take-off at 1945 CST and the Proteus at ~1900 CST. The aircraft performed 1000 ft/min ascents over the CART site, with the Proteus stopping at 43kft to perform level legs parallel to the wind, and the DC-8 performing a level pattern for ~20 minutes at 41kft. At this point, the LASE was seeing a thick low cloud deck over the CART site, and the GSFC SRL was unable to penetrate the clouds from below. The DC-8 aborted its planned flight schedule to perform COAST maneuvers for ~40 minutes. Again, the LASE and CART site conditions were assessed, and the DC-8 mission was aborted altogether, with a flight duration of 3 hours. The Proteus continued with a slow ascent to 54kft followed by a slow descent and return to base.
Despite the uncooperative weather, some useful data was collected for AFWEX, including colocated Vaisala and chilled mirror sonde, LASE, and DC-8 in-situ measurements of water vapor profiles above the clouds colocated with radiation measurements from NASTI, NASTM, FIRSC, and SHIS. Several secondary and non-AFWEX objectives were met, including a ~1 minute over-flight of the "North/South Road" on the approach to the CART site (for land emissivity studies) and COAST data collected at 15kft.

At mid-day, partly cloudy conditions across most of Oklahoma. Low and high clouds over most of Texas and New Mexico were moving south/southeast and a wave from our northwest bringing mid and upper level clouds. Expecting conditions to be better early rather than later. Into the evening, low clouds to our south moved northward but remained mostly south of the CART site. Shortly after sunset, low clouds formed over much of northern Oklahoma. At this point, the interpretation of infrared satellite imagery was complicated by an isothermal boundary layer and an inversion above the cloud tops (see dpcryo5_3.jpg in this directory). The CART site remained under low thick cloud cover until shortly after midnight.

Instrument Status/Comments
CART site:
   CART Raman Lidar Normal operation.
   GSFC Raman Lidar Operated from ~2000 to ~2215 CST.
   MPI DIAL Working on system after flash lamp change.
   WFF Chilled Mirror Sondes Launch at 1920 CST.
   Other 1730 and 2030 CST Vaisala sondes launched at 1830 and 1930 CST.
   LASE Normal operations.
   DLH Normal operations.
   Cryo frost-point/chilled-mirror Normal operations.
   SHIS Normal operations.
   COAST Normal operations. Performed measurements at 3 speeds at 15kft.
   AeroSapient Normal operations.
   NAST-I Normal operations.
   NAST-M Normal operations.
   FIRSC Normal operations.

Dave Tobin, University of Wisconsin